Michael is 30 years older than his daughter. In five years' time, Michael will be 4 times as old as his daughter. How old is Michael now?
Answer by richwmiller(17219)

m=d+30
d=m-30
m+5=4*(d+5)
m+5=4(m-30) +20
m+5=4m-120+20
m+5=4m-100
105=3m
35=m
d=5
check
40=4*10
ok